OpenStudy (anonymous):

quotient rule: (-8)^20/(-8)^18

jimthompson5910 (jim_thompson5910):

When dividing like this, you subtract the corresponding exponents

jimthompson5910 (jim_thompson5910):

Remember if you get a negative result, then you'll have to use the idea that x^(-k) = 1/(x^k) where k is some positive number

OpenStudy (anonymous):

=-8^20-18 =-8^2 =64

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

In this case when terms with the same bases are divided, their exponents are taken away.